explain is the difference between a logarithm of a product and the product of logarithm and give example of each.

Answer by Theo(13342)

logarithm of a product would be:

log(x*y)

product of a logarithm would be:

log(x)*log(y), or it would more likely be:

x * log(y)

I've seen x * log(y) before, but haven't seen log(x) * log(y).

That doesn't mean it doesn't exist. It just means that I haven't seen it.

I have seen log(x) / log(y) before. That's more common.

Some examples:

EXAMPLE OF LOGARITHM OF A PRODUCT

equation is y = 2.773 * 3.402

take log of both sides to get:

log(y) = log(2.773 * 3.402)

this is equivalent to:

log(y) = log(2.773) + log(3.402)

solve for log(y) to get:

log(y) = .974684179

solve for y to get:

y = 9.433746

multiply 2.773 * 3.402 to get:

y = 9.433746

Answers are the same as they should be.

EXAMPLE OF PRODUCT OF A LOGARITHM

y = 7.434^(2.3)

take log of both sides to get:

log(y) = log(7.434^(2.3)

this becomes:

log(y) = 2.3 * log(7.434)

solve for log(y) to get:

log(y) = 2.003811881

solve for y to get:

y = 100.8815812

solve y = 7.434^(2.3) directly using your calculator to get:

y = 100.8815812

Answers are the same as they should be.

EXAMPLE OF DIVISION OF A LOGARITHM

5000 = 2^x

take log of both sides to get:

log(5000) = log(2^x)

this becomes:

log(5000) = x * log(2)

divide both sides by log(2) to get

x = log(5000)/log(2)

solve for x to get:

x = 12.28771238

plug x into original equation to get:

5000 = 2^12.28771238 = 5000